# Comparative Study of State-based Neural Networks for Virtual Analog Audio Effects Modeling

### v5

@@ 3.3 Audio Effects and Datasets
In this section, we introduce the audio effects selected for the study and the composition of their datasets. We proceeded to record the dataset using effects that were at our disposal. In particular, we selected the following effects: overdrive, saturation, equalization, low-pass filter, and compression. Table 2 summarizes all the datasets included in the study. Delay-based effects require a different approach than a direct black box  (Wright & Välimäki,, 2021),  (Mikkonen et al.,, 2023). For this reason, they are not included in this study.
The dataset we use is derived from the following hardware devices: Behringer OD300 overdrive pedal  [^2] https://www.behringer.com/product.html?modelCode=P0608, Behringer Neutron  [^3] https://www.behringer.com/product.html?modelCode=P0CM5’s overdrive module, Behringer Neutron’s filter module in low-pass mode, TC Electronic Bucket Brigade Analog Delay used as a saturator  [^4] https://www.tcelectronic.com/product.html?modelCode=P0EBV, CL 1B TubeTech  [^5] http://www.tube-tech.com/cl-1b-opto-compressor/and Teletronix LA-2A  [^6] https://www.uaudio.com/hardware/la-2a.html optical compressors. We have also collected data from two software VA plugins using  (Fasciani et al.,, 2024b): the Helper Saturator  [^7] https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ  [^8] https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
@@
Similarly to the overdrive, the Behringer Neutron integrates a filter module, which features frequency cutoff and resonance knobs. The filter can be set as low-pass, high-pass, and band-pass. We isolated the module and recorded the dataset at $10$ equally spaced values for the cutoff and $5$ for the resonance, selecting the low-pass filter type. The filter was disconnected from the ADSR envelope module.
@@
Helper Saturator software emulates analog saturation, in particular tube and tape saturation. The software presents four parameters: lowpass and highpass cutoffs, saturation type switch (tape or tube), and saturation level. We recorded the dataset for $10$ different values of the saturation level. The saturation switch was set to tape mode. Lowpass and highpass cutoffs were set to $20$ and $20000$ Hz, respectively.
@@
The software plug-in Universal Audio Pultec Passive Equalizer emulates the tube-based equalizer Pultec EQP-1A. The effect presents low-frequency ([20, 30, 60, 100] Hz) and high-frequency ([3, 4, 5, 8, 10, 12, 16]kHz) switches to select the frequencies to boost or cut using the respective boost and attenuator knobs ([0,10]). Additionally, the emulation also has a bandwidth knob ([0,10]). The equalizer boosts and cuts the selected frequency bands, giving some saturation simultaneously. In this case, we recorded the dataset varying the boost and attenuator knobs at $5$ different equally-spaced values up to half of their range, setting $60$ Hz as the low-frequency switch value and 00 (sharp) as the bandwidth.
@@
The CL 1B and the LA-2A are analog optical compressors. In optical compressors, a lighting-emitting element is fed with the audio signal that illuminates a light-sensitive resistor. The input signal’s amplitude determines the element’s brightness, which, in turn, changes the resistance in the gain attenuation circuit. We recorded the dataset directly from the compressor. The CL 1B compressor dataset presents four parameters: attack, release, ratio, and threshold. We consider three equally spaced values for the ratio, attack, and release time and four values for the threshold. On the other hand, LA-2A presents the peak reduction knob indicating the amount of compression ([00,$100$]) and a mode switch (limiter, compression). We consider 10 equally spaced values for the peak reduction.
@@
LA-2A [^1] Recorded from hardware devices.  [^3] Publicly available. Optical Peak Reduction [0,100] $20$
@@ 3.4 Experimenting and Learning
The models are trained for $200$ epochs and using the Adam (Kingma & Ba,, 2014) optimizer with a gradient norm scaling of $1$  (Pascanu et al.,, 2013). The training was stopped earlier in case of no reduction of validation loss for $10$ epochs. We design a learning rate with an exponential decay according to the formula: $lr=LR*0.25^{e}$, where LR is the initial learning rate and it is set to $3\cdot 10^{-4}$, and $e$ is the number of the epochs. Test losses and evaluation metrics are computed using the model’s weights that minimize the validation loss throughout the training epochs. Finally, the input signal is split into segments of $2400$ samples (equivalent to $50$ ms) to be processed before updating the weights. Considering the network architecture described in Section 3.1, in which a buffer of the $64$ most recent samples is used as input, training layers with states necessitates sequential input-output examples. Consequently, our implementation adopts an input shape of ($B$, $L$, $F$), which is set to ($B$, $1$, $64$), where $B$ denotes the arbitrary batch size, $L$ the sequence length, and $F$ the input dimensionality, often identified as the number of input features. This configuration considers the $64$ input samples as the vector of input features at the current sampling instant.
@@
We adopted the same learning schedule and minibatch size for all models across all datasets. As explained in Section 3.3, we collected the datasets using the same approach and the same input file to minimize the difference among them. However, it is important to highlight that each dataset and modeling task may benefit from specific adjustments to achieve more accurate results. Different datasets and tasks may require slightly different learning rate schedules; an excessively slow or fast learning rate can lead to suboptimal solutions. The number of samples to process before updating the weights referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
The composition of the training, validation, and test sets is identical across datasets. We used $80\%$ for the training set, $10\%$ for the validation set, and the remaining $10\%$ for the test set. For effects with variable control parameters, each $45$s recording in the dataset is associated with a specific combination of the effect parameters. The $80-10-10\%$ split was carried out at the individual recording level, ensuring that all control parameter combinations present an equal share in each subset and are equally used for training and evaluation of the model. The initial $10\%$ of all the recordings are used as the test set. The following $10\%$ portion of the recordings is used as a validation set, while the remaining parts were used as the training set. Minor manual adjustments are made to ensure that splitting points fall within segments of silence.
@@ 4 Results
The evaluation is based on the metrics described in Section 3.2 and presented in Table 4. The MSE loss and the epoch-related epoch resulting in lower validation loss during training are reported in Table 3. The losses and the metrics consider all conditioning scenarios and provide an average error across all combinations of parameters. To further verify the model’s capacity to learn the conditioned effect, we provide a series of plots that compare the target signal with the predictions generated by the four distinct architectures. For these plots, we chose the modeling scenario that we believe to be the most challenging based on the effect parameters used as conditioning factors for the networks. Except for the low-pass filter, this typically involves parameters at the maximum of the selected range since the output signal, under these conditions, is likely to deviate most significantly from the input. Regarding the low-pass filter, the parameters are configured for minimum cutoff and maximum resonance. The plots show the RMSE in Figure  4, and the spectrograms in Figure  5. For datasets including conditioning parameters, the plots showing predictions versus the target are generated for a $2.5$ seconds input signal containing bass, guitar, and drum loop sound. These input signals are part of the respective test sets. To ensure a clearer comparative visualization of the spectral flux, we use only the first $1.25$ seconds.
@@
Table 3: MSE test loss computed on the test set. The test sets include all conditioning values, if applicable. The last column reports the epoch at which the training met the early stopping condition.
Dataset Model MSE Epoch Dataset Model MSE Epoch
OD300 LSTM $\boldsymbol{2.46\cdot 10^{-4}}$ $144$ Neutron LSTM $\boldsymbol{2.27\cdot 10^{-3}}$ $4$
Overdrive ED $3.57\cdot 10^{-4}$ $176$ Overdrive ED $2.79\cdot 10^{-3}$ $3$
LRU $2.98\cdot 10^{-4}$ $170$ LRU $2.43\cdot 10^{-3}$ $100$
S4D $3.23\cdot 10^{-4}$ $165$ S4D $2.62\cdot 10^{-3}$ $8$
S6 $3.39\cdot 10^{-4}$ $172$ S6 $2.41\cdot 10^{-3}$ $28$
Helper LSTM $4.56\cdot 10^{-4}$ $165$ Pultec LSTM $\boldsymbol{7.41\cdot 10^{-4}}$ $19$
Saturator ED $4.41\cdot 10^{-4}$ $138$ Passive ED $1.11\cdot 10^{-3}$ $101$
LRU $\boldsymbol{4.14\cdot 10^{-4}}$ $176$ Equalizer LRU $1.03\cdot 10^{-3}$ $150$
S4D $4.40\cdot 10^{-4}$ $149$ S4D $1.06\cdot 10^{-3}$ $11$
S6 $4.53\cdot 10^{-4}$ $188$ S6 $9.68\cdot 10^{-4}$ $8$
LA-2A LSTM $1.11\cdot 10^{-4}$ $59$ CL 1B LSTM $1.78\cdot 10^{-4}$ $74$
Optical ED $9.54\cdot 10^{-5}$ $90$ Optical ED $\boldsymbol{8.96\cdot 10^{-5}}$ $38$
Compressor LRU $1.14\cdot 10^{-4}$ $146$ Compressor LRU $6.76\cdot 10^{-4}$ $168$
S4D $8.43\cdot 10^{-5}$ $62$ S4D $1.22\cdot 10^{-4}$ $14$
S6 $\boldsymbol{8.25\cdot 10^{-5}}$ $42$ S6 $1.01\cdot 10^{-4}$ $19$
Neutron LSTM $4.07\cdot 10^{-3}$ $46$
Filter ED $3.36\cdot 10^{-3}$ $9$
LRU $\boldsymbol{2.43\cdot 10^{-3}}$ $188$
S4D $2.75\cdot 10^{-3}$ $12$
S6 $3.16\cdot 10^{-3}$ $7$
Table 4: ESR, $M_{RMSE}$, $M_{SF}$, and $M_{STFT}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
Dataset Model ESR $M_{RMSE}$ $M_{SF}$ $M_{STFT}$ $M_{MFCC}$
OD300 LSTM $\boldsymbol{1.06\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $1.60\cdot 10^{-2}$ $3.56\cdot 10^{-1}$ $6.08\cdot 10^{-1}$
Overdrive ED $1.54\cdot 10^{-1}$ $1.30\cdot 10^{-2}$ $1.19\cdot 10^{-2}$ $3.52\cdot 10^{-1}$ $5.80\cdot 10^{-1}$
LRU $2.35\cdot 10^{-1}$ $1.65\cdot 10^{-2}$ $1.77\cdot 10^{-2}$ $4.55\cdot 10^{-1}$ $1.05$
S4D $1.40\cdot 10^{-1}$ $\boldsymbol{1.20\cdot 10^{-2}}$ $\boldsymbol{8.50\cdot 10^{-3}}$ $3.11\cdot 10^{-1}$ $\boldsymbol{5.38\cdot 10^{-1}}$
S6 $1.46\cdot 10^{-1}$ $1.25\cdot 10^{-2}$ $1.04\cdot 10^{-2}$ $\boldsymbol{2.78\cdot 10^{-1}}$ $5.61\cdot 10^{-1}$
Neutron LSTM $\boldsymbol{2.50\cdot 10^{-1}}$ $\boldsymbol{1.86\cdot 10^{-2}}$ $5.74\cdot 10^{-3}$ $4.80\cdot 10^{-1}$ $3.49\cdot 10^{-1}$
Overdrive ED $3.07\cdot 10^{-1}$ $2.36\cdot 10^{-2}$ $6.33\cdot 10^{-3}$ $5.64\cdot 10^{-1}$ $3.80\cdot 10^{-1}$
LRU $2.70\cdot 10^{-1}$ $1.94\cdot 10^{-2}$ $6.64\cdot 10^{-3}$ $\boldsymbol{3.40\cdot 10^{-1}}$ $3.39\cdot 10^{-1}$
S4D $2.88\cdot 10^{-1}$ $2.21\cdot 10^{-2}$ $5.71\cdot 10^{-3}$ $5.20\cdot 10^{-1}$ $3.59\cdot 10^{-1}$
S6 $2.65\cdot 10^{-1}$ $1.89\cdot 10^{-2}$ $\boldsymbol{4.97\cdot 10^{-3}}$ $3.86\cdot 10^{-1}$ $\boldsymbol{3.05\cdot 10^{-1}}$
Helper LSTM $2.67\cdot 10^{-1}$ $1.05\cdot 10^{-2}$ $2.78\cdot 10^{-2}$ $5.97\cdot 10^{-1}$ $4.72\cdot 10^{-1}$
Saturator ED $2.58\cdot 10^{-1}$ $1.06\cdot 10^{-2}$ $1.29\cdot 10^{-2}$ $5.33\cdot 10^{-1}$ $4.14\cdot 10^{-1}$
LRU $3.46\cdot 10^{-1}$ $\boldsymbol{9.72\cdot 10^{-3}}$ $1.42\cdot 10^{-2}$ $5.44\cdot 10^{-1}$ $\boldsymbol{2.46\cdot 10^{-1}}$
S4D $\boldsymbol{2.57\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $\boldsymbol{1.22\cdot 10^{-2}}$ $5.41\cdot 10^{-1}$ $3.67\cdot 10^{-1}$
S6 $2.65\cdot 10^{-1}$ $1.01\cdot 10^{-2}$ $1.31\cdot 10^{-2}$ $\boldsymbol{4.28\cdot 10^{-1}}$ $3.88\cdot 10^{-1}$
Pultec LSTM $\boldsymbol{2.15\cdot 10^{-1}}$ $1.29\cdot 10^{-2}$ $8.85\cdot 10^{-3}$ $9.33\cdot 10^{-1}$ $2.06\cdot 10^{-1}$
Passive ED $3.22\cdot 10^{-1}$ $1.24\cdot 10^{-2}$ $2.49\cdot 10^{-3}$ $4.94\cdot 10^{-1}$ $1.91\cdot 10^{-1}$
Equalizer LRU $3.00\cdot 10^{-1}$ $\boldsymbol{1.16\cdot 10^{-2}}$ $1.99\cdot 10^{-3}$ $\boldsymbol{1.73\cdot 10^{-1}}$ $9.81\cdot 10^{-2}$
S4D $3.08\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $\boldsymbol{1.79\cdot 10^{-3}}$ $4.09\cdot 10^{-1}$ $\boldsymbol{9.44\cdot 10^{-2}}$
S6 $2.80\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $3.89\cdot 10^{-3}$ $2.87\cdot 10^{-1}$ $1.26\cdot 10^{-1}$
LA-2A LSTM $1.03\cdot 10^{-1}$ $4.51\cdot 10^{-3}$ $1.12\cdot 10^{-2}$ $4.37\cdot 10^{-1}$ $3.10\cdot 10^{-1}$
Optical ED $8.91\cdot 10^{-2}$ $4.10\cdot 10^{-3}$ $7.28\cdot 10^{-3}$ $2.30\cdot 10^{-1}$ $2.30\cdot 10^{-1}$
Compressor LRU $1.06\cdot 10^{-1}$ $4.67\cdot 10^{-3}$ $7.84\cdot 10^{-3}$ $2.70\cdot 10^{-1}$ $2.93\cdot 10^{-1}$
S4D $7.87\cdot 10^{-2}$ $\boldsymbol{3.51\cdot 10^{-3}}$ $1.00\cdot 10^{-2}$ $2.52\cdot 10^{-1}$ $2.53\cdot 10^{-1}$
S6 $\boldsymbol{7.71\cdot 10^{-2}}$ $3.63\cdot 10^{-3}$ $\boldsymbol{6.26\cdot 10^{-3}}$ $\boldsymbol{1.84\cdot 10^{-1}}$ $\boldsymbol{2.12\cdot 10^{-1}}$
CL 1B LSTM $2.92\cdot 10^{-1}$ $9.33\cdot 10^{-3}$ $5.21\cdot 10^{-3}$ $3.56\cdot 10^{-1}$ $2.26\cdot 10^{-1}$
Optical ED $\boldsymbol{1.46\cdot 10^{-1}}$ $\boldsymbol{4.05\cdot 10^{-3}}$ $\boldsymbol{4.18\cdot 10^{-3}}$ $4.17\cdot 10^{-1}$ $\boldsymbol{2.09\cdot 10^{-1}}$
Compressor LRU $1.88\cdot 10^{-2}$ $6.38\cdot 10^{-3}$ $5.49\cdot 10^{-3}$ $4.38\cdot 10^{-1}$ $4.52\cdot 10^{-1}$
S4D $2.00\cdot 10^{-1}$ $5.02\cdot 10^{-3}$ $6.12\cdot 10^{-3}$ $\boldsymbol{2.79\cdot 10^{-1}}$ $2.93\cdot 10^{-1}$
S6 $1.67\cdot 10^{-1}$ $4.41\cdot 10^{-3}$ $6.23\cdot 10^{-3}$ $4.99\cdot 10^{-1}$ $2.36\cdot 10^{-1}$
Neutron LSTM $1.06$ $3.15\cdot 10^{-2}$ $3.03\cdot 10^{-2}$ $9.33\cdot 10^{-1}$ $6.57\cdot 10^{-1}$
Filter ED $8.77\cdot 10^{-1}$ $2.18\cdot 10^{-2}$ $2.21\cdot 10^{-2}$ $9.38\cdot 10^{-1}$ $6.32\cdot 10^{-1}$
LRU $\boldsymbol{6.58\cdot 10^{-1}}$ $\boldsymbol{1.66\cdot 10^{-2}}$ $\boldsymbol{1.95\cdot 10^{-2}}$ $\boldsymbol{7.51\cdot 10^{-1}}$ $\boldsymbol{5.98\cdot 10^{-1}}$
S4D $7.18\cdot 10^{-1}$ $1.97\cdot 10^{-2}$ $2.00\cdot 10^{-2}$ $9.01\cdot 10^{-1}$ $6.28\cdot 10^{-1}$
S6 $8.23\cdot 10^{-1}$ $2.17\cdot 10^{-2}$ $2.18\cdot 10^{-2}$ $8.94\cdot 10^{-1}$ $6.27\cdot 10^{-1}$

### v6

@@ 3.3 Audio Effects and Datasets
In this section, we introduce the audio effects selected for the study and the composition of their datasets. We proceeded to record the dataset using effects that were at our disposal. In particular, we selected the following effects: overdrive, saturation, equalization, low-pass filter, and compression. Table 2 summarizes all the datasets included in the study. Delay-based effects require a different approach than a direct black box  Wright & Välimäki, (2021),  Mikkonen et al., (2023). For this reason, they are not included in this study.
The dataset we use is derived from the following hardware devices: Behringer OD300 overdrive pedal  [^2] https://www.behringer.com/product.html?modelCode=P0608, Behringer Neutron  [^3] https://www.behringer.com/product.html?modelCode=P0CM5’s overdrive module, Behringer Neutron’s filter module in low-pass mode, TC Electronic Bucket Brigade Analog Delay used as a saturator  [^4] https://www.tcelectronic.com/product.html?modelCode=P0EBV, CL 1B TubeTech  [^5] http://www.tube-tech.com/cl-1b-opto-compressor/and Teletronix LA-2A  [^6] https://www.uaudio.com/hardware/la-2a.html optical compressors. We have also collected data from two software VA plugins using  Fasciani et al., (2024b): the Helper Saturator  [^7] https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ  [^8] https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
@@
Similarly to the overdrive, the Behringer Neutron integrates a filter module, which features frequency cutoff and resonance knobs. The filter can be set as low-pass, high-pass, and band-pass. We isolated the module and recorded the dataset at $10$ equally spaced values for the cutoff and $5$ for the resonance, selecting the low-pass filter type. The filter was separated from the envelope module, which can be used to modulate it.
@@
The Helper Saturator software emulates analog saturation, in particular tube and tape saturation. The software presents four parameters: lowpass and highpass cutoffs, saturation type switch (tape or tube), and saturation level. We recorded the dataset for $10$ different values of the saturation level. The saturation switch was set to tape mode. Lowpass and highpass cutoffs were set to $20$ and $20000$ Hz, respectively.
@@
The software plug-in Universal Audio Pultec Passive Equalizer emulates the tube-based equalizer Pultec EQP-1A. The effect presents low-frequency ([20, 30, 60, 100] Hz) and high-frequency ([3, 4, 5, 8, 10, 12, 16]kHz) switches to select the frequencies to boost or cut using the respective boost and attenuator knobs ([0,10]). Additionally, the emulation also has a bandwidth knob ([0,10]). The equalizer boosts and cuts the selected frequency bands, giving some saturation simultaneously. In this case, we recorded the dataset varying the boost and attenuator knobs at $5$ different equally-spaced values up to half of their range, setting $60$ Hz as the low-frequency switch value and 0 (sharp) as the bandwidth.
@@
The CL 1B and the LA-2A are analog optical compressors. In optical compressors, a lighting-emitting element is fed with the audio signal that illuminates a light-sensitive resistor. The input signal’s amplitude determines the element’s brightness, which, in turn, changes the resistance in the gain attenuation circuit. We recorded the dataset directly from the compressor. The CL 1B compressor dataset presents four parameters: attack, release, ratio, and threshold. We consider three equally spaced values for the ratio, attack, and release time and four values for the threshold. On the other hand, LA-2A presents the peak reduction knob indicating the amount of compression ([0,$100$]) and a mode switch (limiter, compression). We consider $10$ equally spaced values for the peak reduction.
@@
LA-2A [^1] Recorded from hardware devices. Optical Peak Reduction [0,100] 2020
@@ 3.4 Experimenting and Learning
The models are trained for $200$ epochs and use the Adam Kingma & Ba, (2014) optimizer with a gradient norm clipped at 11  Pascanu et al., (2013). The training was stopped earlier in case of no reduction of validation loss for $10$ epochs. We design a learning rate with an exponential decay according to the formula: $lr=LR*0.25^{e}$, where LR is the initial learning rate and it is set to $3\cdot 10^{-4}$, and $e$ is the number of epochs. Test losses and evaluation metrics are computed using the model’s weights that minimize the validation loss throughout the training epochs. Finally, the input signal is split into segments of $2400$ samples (equivalent to $50$ ms) to be processed before updating the weights. Considering the network architecture described in Section 3.1. Consequently, our implementation adopts an input shape of ($B$, $L$, $F$), which is set to ($B$, $2400$, $64$), where $B$ denotes the arbitrary batch size, $L$ the sequence length, and $F$ the input dimensionality, often identified as the number of input features. This configuration is used within the framework of Truncated Backpropagation Through Time (TBPTT).
@@
We adopted the same learning schedule and minibatch size for all models across all datasets. As explained in Section 3.3, we collected the datasets using the same approach and the same input file to minimize the differences among them. However, it is important to highlight that each dataset and modeling task may benefit from specific adjustments to achieve more accurate results. Different datasets and tasks may require slightly different learning rate schedules; an excessively slow or fast learning rate can lead to suboptimal solutions. The number of samples to process before updating the weights, referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
To ensure robust evaluation of the models, we divided the collected data into $5$ distinct compositions of training, validation, and test sets. This strategy aimed to cross-validate the models across different data splits. In each composition, $80\%$ of the data is allocated for training, $10\%$ is reserved for validation, and the remaining $10\%$ constitutes the test set. Importantly, the validation and test sets in each composition include different data signals, ensuring that the model never uses the same audio signals in both validation and testing processes. This approach prevents any bias towards certain data signals and allows for a more accurate assessment of the model’s performance across varying conditions and inputs.
For audio effects with variable control parameters, each $45$ s recording in the dataset corresponds to a specific combination of effect parameters. To maintain consistency across training, validation, and testing, we employ a $80-10-10\%$ split at the level of individual recordings. This ensures that all combinations of control parameters are proportionally represented in each subset, providing balanced usage for both training the model and evaluating its performance. Additionally, minor manual adjustments were made to the splitting points to ensure they occur during silent segments, preventing abrupt changes in audio dynamics that could affect model accuracy.
The statistical significance of the results is analyzed using the Friedman test and the Wilcoxon signed-rank test  demvsar2006statistical; rainio2024evaluation. Both are non-parametric statistical tests employed to determine whether there are statistically significant differences between the medians of samples. These tests are well-suited for data that does not meet the assumptions required for parametric tests, such as normality.
In this case, where we are evaluating $5$ models across $5$ different dataset compositions, the Friedman test is the most appropriate for assessing statistical significance. However, due to the limited number of dataset compositions, the reliability of the Friedman test might be reduced. To address this, we also conduct separate Wilcoxon signed-rank tests for each model pair. While this test is useful for assessing significant differences between two models, it’s important to note that it is not ideal for comparing multiple models. Specifically, conducting multiple Wilcoxon tests increases the risk of incorrectly finding significant differences due to the numerous comparisons involved.
@@ 4 Results
The MSE loss, the metrics described in Section 3.2, and the mean of the epochs resulting in lower validation loss during training are reported in Table 3. The losses and the metrics consider all conditioning scenarios and provide an average error across all combinations of parameters. The average epoch at which the training met the early stopping condition is also reported in the table. LRU is generally the slowest to converge, taking significantly more epochs to reach the same range of errors as the other models. On the other hand, other variants are dataset-specific, with a tendency for the SSM-based model, such as S4D and S6, to converge faster and with a similar number of epochs. Similarly, both LSTM and ED require a comparable number of epochs, although they generally converge more slowly than models based on SSM. This slower convergence is attributed to their similar layer formulation. On the other hand, LRU utilizes specific initializations and normalization schemes, distinguishing it from traditional recurrent and SSM architectures.
@@
Table 3: Mean values of MSE loss, ESR, $M_{NRMSE}$, $M_{SF}$, and $M_{STFT}$ metrics computed across different test sets, with each set linked to a model trained on a distinct dataset split. The last column reports the average epoch at which the training met the early stopping condition.
Dataset Model MSE ESR $M_{NRMSE}$ $M_{SF}$ $M_{STFT}$ Epoch
OD300 LSTM $\boldsymbol{2.92\cdot 10^{-4}}$ $\boldsymbol{1.13\cdot 10^{-1}}$ $\boldsymbol{2.69\cdot 10^{-1}}$ $\boldsymbol{4.56\cdot 10^{-1}}$ $7.50\cdot 10^{-1}$ $157$
Overdrive ED $3.59\cdot 10^{-4}$ $2.00\cdot 10^{-1}$ $2.90\cdot 10^{-1}$ $5.66\cdot 10^{-1}$ $7.78\cdot 10^{-1}$ $158$
LRU $4.57\cdot 10^{-4}$ $1.77\cdot 10^{-1}$ $3.44\cdot 10^{-1}$ $5.35\cdot 10^{-1}$ $8.12\cdot 10^{-1}$ $170$
S4D $3.14\cdot 10^{-4}$ $1.22\cdot 10^{-1}$ $2.88\cdot 10^{-1}$ $4.89\cdot 10^{-1}$ $\boldsymbol{7.40\cdot 10^{-1}}$ $189$
S6 $4.45\cdot 10^{-4}$ $1.72\cdot 10^{-1}$ $3.57\cdot 10^{-1}$ $5.43\cdot 10^{-1}$ $8.82\cdot 10^{-1}$ $153$
Neutron LSTM $\boldsymbol{5.64\cdot 10^{-4}}$ $\boldsymbol{6.21\cdot 10^{-2}}$ $\boldsymbol{2.51\cdot 10^{-1}}$ $\boldsymbol{3.25\cdot 10^{-1}}$ $5.45\cdot 10^{-1}$ $18$
Overdrive ED $7.84\cdot 10^{-4}$ $8.64\cdot 10^{-2}$ $2.87\cdot 10^{-1}$ $3.53\cdot 10^{-1}$ $\boldsymbol{5.42\cdot 10^{-1}}$ $54$
LRU $9.65\cdot 10^{-4}$ $1.10\cdot 10^{-1}$ $3.10\cdot 10^{-1}$ $4.09\cdot 10^{-1}$ $5.98\cdot 10^{-1}$ $106$
S4D $1.12\cdot 10^{-3}$ $1.30\cdot 10^{-1}$ $4.20\cdot 10^{-1}$ $5.33\cdot 10^{-1}$ $5.98\cdot 10^{-1}$ $10$
S6 $1.16\cdot 10^{-3}$ $1.32\cdot 10^{-1}$ $4.42\cdot 10^{-1}$ $5.08\cdot 10^{-1}$ $6.58\cdot 10^{-1}$ $10$
Helper LSTM $3.47\cdot 10^{-4}$ $1.91\cdot 10^{-1}$ $3.81\cdot 10^{-1}$ $\boldsymbol{4.65\cdot 10^{-1}}$ $6.88\cdot 10^{-1}$ $180$
Saturator ED $3.49\cdot 10^{-4}$ $1.66\cdot 10^{-1}$ $3.82\cdot 10^{-1}$ $4.82\cdot 10^{-1}$ $6.70\cdot 10^{-1}$ $94$
LRU $6.11\cdot 10^{-4}$ $3.36\cdot 10^{-1}$ $4.96\cdot 10^{-1}$ $5.54\cdot 10^{-1}$ $8.90\cdot 10^{-1}$ $157$
S4D $\boldsymbol{2.93\cdot 10^{-4}}$ $\boldsymbol{1.62\cdot 10^{-1}}$ $\boldsymbol{3.54\cdot 10^{-3}}$ $4.72\cdot 10^{-1}$ $6.59\cdot 10^{-1}$ $109$
S6 $2.99\cdot 10^{-4}$ $1.63\cdot 10^{-1}$ $3.57\cdot 10^{-1}$ $4.85\cdot 10^{-1}$ $\boldsymbol{6.57\cdot 10^{-1}}$ $176$
Pultec LSTM $\boldsymbol{4.26\cdot 10^{-5}}$ $\boldsymbol{1.14\cdot 10^{-2}}$ $\boldsymbol{9.63\cdot 10^{-3}}$ $\boldsymbol{7.11\cdot 10^{-2}}$ $\boldsymbol{1.85\cdot 10^{-1}}$ $66$
Passive ED $1.43\cdot 10^{-4}$ $3.49\cdot 10^{-2}$ $56\cdot 10^{-1} 1.1$$18\cdot 10^{-1}$ $2.92\cdot 10^{-1}$ $130$
Equalizer LRU $2.19\cdot 10^{-4}$ $5.60\cdot 10^{-2}$ $1.98\cdot 10^{-1}$ $2.17\cdot 10^{-1}$ $3.62\cdot 10^{-1}$ $173$
S4D $1.62\cdot 10^{-4}$ $6.97\cdot 10^{-2}$ $8.73\cdot 10^{-2}$ $7.46\cdot 10^{-2}$ $2.28\cdot 10^{-1}$ 88
S6 $4.79\cdot 10^{-5}$ $1.20\cdot 10^{-2}$ $1.00\cdot 10^{-1}$ $8.41\cdot 10^{-1}$ $1.96\cdot 10^{-1}$ $56$
LA-2A LSTM $90\cdot 10^{-4} 1.1$$11\cdot 10^{-1}$ $11\cdot 10^{-1} 3.3$$34\cdot 10^{-1}$ $6.61\cdot 10^{-1}$ $57$
Optical ED $4.22\cdot 10^{-5}$ $2.36\cdot 10^{-2}$ $1.60\cdot 10^{-1}$ $1.78\cdot 10^{-1}$ $3.47\cdot 10^{-1}$ $70$
Comp LRU $5.08\cdot 10^{-5}$ $2.82\cdot 10^{-2}$ $1.67\cdot 10^{-1}$ $2.22\cdot 10^{-1}$ $3.97\cdot 10^{-1}$ $154$
S4D $3.76\cdot 10^{-5}$ $2.07\cdot 10^{-2}$ $1.47\cdot 10^{-1}$ $1.62\cdot 10^{-1}$ $3.44\cdot 10^{-1}$ $54$
S6 $\boldsymbol{3.60\cdot 10^{-5}}$ $\boldsymbol{1.96\cdot 10^{-2}}$ $\boldsymbol{1.43\cdot 10^{-1}}$ $\boldsymbol{1.53\cdot 10^{-1}}$ $\boldsymbol{3.31\cdot 10^{-1}}$ $22$
CL 1B LSTM $3.37\cdot 10^{-4}$ $2.72\cdot 10^{-2}$ $4.88\cdot 10^{-1}$ $5.97\cdot 10^{-1}$ $9.16\cdot 10^{-1}$ $35$
Optical ED $\boldsymbol{2.68\cdot 10^{-5}}$ $\boldsymbol{2.01\cdot 10^{-2}}$ $\boldsymbol{1.20\cdot 10^{-1}}$ $\boldsymbol{1.43\cdot 10^{-1}}$ $\boldsymbol{3.51\cdot 10^{-1}}$ $52$
Comp LRU $6.21\cdot 10^{-5}$ $5.02\cdot 10^{-2}$ $1.96\cdot 10^{-1}$ $2.23\cdot 10^{-1}$ $4.59\cdot 10^{-1}$ $168$
S4D $1.09\cdot 10^{-4}$ $9.06\cdot 10^{-2}$ $2.78\cdot 10^{-1}$ $2.95\cdot 10^{-1}$ $5.83\cdot 10^{-1}$ $28$
S6 $6.64\cdot 10^{-5}$ $5.12\cdot 10^{-2}$ $1.83\cdot 10^{-1}$ $2.00\cdot 10^{-1}$ $4.66\cdot 10^{-1}$ $14$
Neutron LSTM $3.18\cdot 10^{-3}$ $7.03\cdot 10^{-1}$ $8.28\cdot 10^{-1}$ $8.62\cdot 10^{-1}$ $1.19$ $76$
Filter ED $2.91\cdot 10^{-3}$ $6.38\cdot 10^{-1}$ $\boldsymbol{6.82\cdot 10^{-1}}$ $\boldsymbol{7.57\cdot 10^{-1}}$ $1.13$ $54$
LRU $2.62\cdot 10^{-3}$ $5.77\cdot 10^{-1}$ $7.18\cdot 10^{-1}$ $\cdot 10^{-1} 1.081.08$ $179$
S4D $\boldsymbol{2.52\cdot 10^{-3}}$ $\boldsymbol{5.56\cdot 10^{-1}}$ $6.94\cdot 10^{-1}$ $7.90\cdot 10^{-1}$ $1.06$ $16$
S6 $2.62\cdot 10^{-3}$ $5.77\cdot 10^{-1}$ $7.21\cdot 10^{-1}$ $8.11\cdot 10^{-1}$ $\boldsymbol{1.04}$ 77